# N2: A Unified Python Package and Test Bench for Nearest Neighbor-Based Matrix Completion

### v1

@@ 2.3 New variant: Auto nearest neighbors
$\mathsf{TSNN}$ is a generalization of $\mathsf{RowNN}$ and $\mathsf{ColNN}$ by setting one of the tuning parameters to zero (see Tab. 1), whereas the idea underlying $\mathsf{DRNN}$ is fundamentally different from that of $\mathsf{TSNN}$; $\mathsf{DRNN}$ debiases a naive combination of $\mathsf{RowNN}$ and $\mathsf{ColNN}$ whereas $\mathsf{TSNN}$ simply boosts the number of measurements averaged upon, thereby gaining from lower variance. So we simply interpolate the two methods for some hyper-parameter $\alpha\in[0,1]$; see Tab. 1. Notably the hyper-parameter $\eta$ for both $\mathsf{DRNN}$ and $\mathsf{TSNN}$ are identical when interpolated.
Suppose $\mu_{i,t}=\theta_{i,t}+\varepsilon_{i,t}$ in 1 where $\varepsilon_{i,t}$ are centered i.i.d. sub-Gaussian distributions across $i$ and $t$. When $\sigma$ is large in magnitude, $\mathsf{TSNN}$ denoises the estimate by averaging over more samples, hence providing a superior performance compared to $\mathsf{DRNN}$ in a noisy scenario. When $\sigma$ is small so that bias of nearest neighbor is more prominent, $\mathsf{DRNN}$ effectively debiases the estimate so as to provide a superior performance compared to $\mathsf{TSNN}$. The linear interpolator $\mathsf{AutoNN}$ automatically adjusts to the underlying noise level and debiases or denoises accordingly; such property is critical when applying nearest neighbors to real world data set where the noise level is unknown. We refer to Fig. 1 for visual evidence.
@@
Figure 1: Error scaling for certain NN variants in synthetic experiments. See Sec. D.1 for details on the data-generating process and how the signal-to-noise ratio (SNR) is defined. Each point corresponds to the mean absolute error ± 1 standard error across 30 trials.
@@ 3 $\mathbf{N}^{2}$ Package and Interface
We now present our unified Python package, $\mathbf{N}^{2}$, for nearest neighbor algorithms for matrix completion. In particular, we provide a class structure which abstracts the estimation procedure utilized in each different nearest neighbor method as well as the the Distance and Average modules described above in Sec. 2. On top of that, our library facilitates easy extensions to other nearest neighbors algorithms and other data types on top of scalars and distributions. For example, as long as a distance and average notion are well defined, our library can be easily applied to a matrix of images or text strings.
@@
EstimationMethod classes contain the logic to impute a missing entry such as how to use calculated distances. In the context of the modules in Sec. 2, this class determines the weighting function or, in the context of $\mathsf{DRNN}$ and $\mathsf{AutoNN}$, how to compose estimates. We separate this from the DataType abstraction because several estimation methods can be used for multiple data types. For example, RowRowEstimator implements the $\mathsf{RowNN}$ procedure for any data type given to it, such as scalars or distributions.
@@
To use our library, a user simply has to instantiate a composite class NearestNeighborImputer with their EstimationMethod and DataType of choice. We provide constructor functions to automatically create popular NearestNeighborImputer classes such as a two-sided nearest neighbor estimator with the scalar data type. From a design pattern point of view, this is known as a Composite design pattern (GHJV, 93, pg. 163). We use this design pattern so that anyone looking to customize the estimation procedure can do so for any kind of data type simultaneously. Similarly, with the exception of doubly robust estimators, each estimation procedure works out of the box with any data type that implements the DataType abstract class. The Doubly robust estimation method does not work out of the box with distributions because a subtraction operation is not well defined in the distribution space.
@@ 4 $\mathbf{N}^{2}$-Bench and Results
In this section, we evaluate several nearest neighbor algorithms provided by our library, $\mathbf{N}^{2}$, on real-world data. As part of our package, we include data loaders which automatically download the necessary datasets and format them for evaluation. These datasets and loaders comprise our proposed benchmark for nearest neighbor matrix completion algorithms, $\mathbf{N}^{2}$-Bench. We also test several existing popular matrix completion techniques (HMLZ (15); Cha (15)). For details on our experimental setup, computing hardware, and boxplot generation, see App. D.
@@ 4.1 Personalized healthcare: HeartSteps
The HeartSteps V1 study (HeartSteps study for short) is a clinical trial designed to measure the efficacy of the HeartSteps mobile application for encouraging non-sedentary activity KSS+ (19). The HeartSteps V1 data and its subsequent extensions have been widely used for benchmarking a variety of tasks including counterfactual inference of treatment effect DTT+22a ; CFC+ (24), reinforcement learning for intervention selection LGKM (20), and micro-randomized trial design QWC+ (22). In the HeartSteps study, $N=37$ participants were under a 6-week period micro-randomized trial, where they were provided with a mobile application and an activity tracker. Participants independently received a notification with probability $p=0.6$ for $5$ pre-determined decision points per day for 40 days ($T=200$). We denote observed entries $Z_{i,t}$ as the mean participant step count for one hour after a notification was sent and unobserved entries as the unknown step count for decision points where no notification was sent. Our task is to estimate the counterfactual outcomes: the participant’s step count should they have received a different treatment (notification or no notification) than they did at specific time points during the study.
@@
In Fig. 2(a), we compare the absolute error of the imputed values across the nearest neighbor and baseline methods. The scalar nearest neighbor methods far out-perform USVT and are on par with SoftImpute. The two distributional nearest neighbor methods far outperform all methods operating in the scalar setting; it suggests that matching by distributions collect more homogeneous neighbors, thereby decreases the bias of the method, compared to matching only the first moments as done in most scalar matrix nearest neighbor methods.
In Fig. 2 panel (b), we show an example of an imputed entry in the distributional nearest neighbors setting. In this case, the ground truth distribution is bimodal, as the participant was largely sedentary (0 steps) with small amounts of activity. While both $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ capture the sedentary behavior of the participant, $\mathsf{KernelNN}$ is able to recover the bimodality of the original distribution whereas $\mathsf{W_{2}NN}$ cannot.
@@ 4.2 Movie recommendations: MovieLens
Figure 3: MovieLens: Estimation error for a random subsample of size 500. For experimental settings and discussion see Sec. 4.2.
The MovieLens 1M dataset HK (15) contains 1 million ratings (1–5 stars) from 6,040 users on 3,952 movies. Collaborative filtering on MovieLens has long been a benchmark for matrix-completion methods: neighborhood-based algorithms SKKR (01), latent-factor models KBV (09), and, more recently, nearest neighbors interpreted as blind regression under a latent–variable model LSSY (19). These assist practitioners in data-driven recommendation systems, since more accurate rating imputation directly drives better personalized suggestions and user engagement. This is a standard scalar matrix completion problem with $N=6,040$ and $T=3,952$. Each rating is an integer in $\{1,\dots,5\}$. The dataset has a very high percentage of missing values: 95.53% missing. Our task is to estimate unobserved ratings using various matrix completion algorithms. We benchmark the performance of nearest neighbors against matrix factorization by measuring absolute error on held-out ratings. See Sec. D.3 for additional details on the dataset.
@@
We fit the nearest neighbor methods using a random sample of size 100 from the first 80% of the dataset to choose nearest neighbor hyperparameters via cross-validation.We then test the method on a random subsample of size 500 from the last 20% of the dataset. As observed in Fig. 3, all nearest neighbor methods have a lower average error than USVT and a much lower standard deviation of errors, with $\mathsf{ColNN}$, $\mathsf{RowNN}$, $\mathsf{DRNN}$, and $\mathsf{AutoNN}$ performing the best out of the nearest neighbor methods. SoftImpute performs on par with the nearest neighbor methods. Note that the nearest neighbor methods perform well even while only being trained on a tiny subset of the data of size 100 out of the 1 million ratings available.
@@ 4.3 Counterfactual inference for panel data: Proposition 99
Figure 4: Nearest neighbor methods generate high-fidelity synthetic controls in counterfactual inference for panel data. For exact settings and further discussion see Sec. 4.3.
Next we consider a panel data setting, where our goal is to estimate the effect of the California Tobacco Tax and Health Protection Act of 1988 (a.k.a. Proposition 99) on annual state-level cigarette consumption111measured as per capita cigarette sales in packs. By definition, the counterfactual cigarette consumption in California—had Proposition 99 never been enacted—is not observed. ADH (10) introduce the notion of a “synthetic control” to serve as a proxy for this unobserved value based on “neighboring” control states that never instituted a tobacco tax. These states are not close in a geographical sense, but rather close due to similarities in other covariates [^2] GDP per capita, beer consumption, percent aged 15–24, and cigarette retail prices. We take a different approach and use only the observed cigarette consumption levels from the control states, of which there are 38 in total. Thus, we frame our problem as a scalar matrix completion problem with $N=39$ and $T=31$ (see 1). The last row in the matrix corresponds to the state of California.
@@
For each method, we use a 64-16-20 train-validation-test split and use cross validation to fit any hyperparameters. Fig. 4 plots the various synthetic controls for California (left) and absolute error of each method on the 38 control states, for which we do observe the no-treatment values (right). From Fig. 4(a), we see that nearest neighbor methods, in particular $\mathsf{TSNN}$ and $\mathsf{RowNN}$, are roughly on par with the gold-standard synthetic control method of ADH (10) (“SC”) for estimating California’s counterfactual cigarette consumption in the post-intervention period (after 1989). This is despite the fact that the nearest neighbor methods rely on less information for the estimation task. From Fig. 4(b), we see that all nearest neighbor methods, with the exception of $\mathsf{ColNN}$, achieve similar error levels as the synthetic control baseline. $\mathsf{RowNN}$ achieves even lower error levels. See supplementary experiment details in Sec. D.4.
@@ 4.4 Efficient LLM evaluation: PromptEval
The rapid advancement of LLMs have placed them at the center of many modern machine learning systems, from chatbots to aids in medical education GHC+ (25). In practice, system architects want to strike the right balance of real-world performance and cost, but navigating this Pareto frontier is a daunting task. 2024 alone saw at least 10 new models from Anthropic, Google, Meta, and OpenAI, not even counting the multitude of open-source fine-tuned models built on top of these. On specific tasks, smaller, fine-tuned models may even outperform the latest frontier models, in addition to being more cost effective.
We investigate how matrix completion, specifically nearest neighbor methods, can alleviate some of these burdens. We use the PromptEval dataset PXW+ (24), which evaluates $15$ open-source language models (ranging in size from 3B to 70B parameters) and $100$ different prompting techniques across the 57 tasks of the MMLU benchmark (HBB+, 20). In practice, the performance of a model depends—sometimes dramatically—on the precise input prompt. This suggests that we need to consider the performance of a model across a wide range of prompts, rather than any one prompt in particular. Thus, we model this problem as a distributional matrix completion problem with $N=15$, $T=57$, and $n=100$. Given one of 57 tasks, we aim to accurately characterize the performance of each model without resorting to exhaustive evaluation. Nearest neighbors leverage commonalities across models and tasks to estimate the performance distribution of each entry, which was otherwise not considered in PXW+ (24); previous literature achieves efficient evaluation per model and task in isolation without leveraging any across model / task information.
@@
Figure 5: Distributional nearest neighbor methods enable efficient LLM evaluation on MMLU. We estimate LLM score distributions across all models and tasks given only a limited number of model-task evaluations, determined by the propensity $p$. See Sec. 4.4 for a detailed discussion.
@@
We randomly include each entry in the matrix independently with probability $p\in\mathopen{}\mathclose{{}\left\{0.3,0.5,0.7}\right\}$ and impute the missing entries using the $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ methods of Tab. 1. For each method, we consider both the the row-wise and column-wise variants. Fig. 5(a) reports the mean Kolmogorov-Smirnov (KS) distance between the imputed and ground-truth distributions across the entries in the test set for varying missingness values. As expected, estimation error decreases as $p$ increases. Fig. 5(b) visualizes the imputed distributions using row-wise $\mathsf{KernelNN}$ and column-wise $\mathsf{W_{2}NN}$ (at $p=0.7$) for a select entry, along with the ground-truth distribution. Even with 30% of matrix entries missing, distributional NN methods are able to recover the underlying distribution.
@@ Appendix A Structural assumptions
[…]
The first assumption specifies the factor structure on the mean; that is, there exists latent factors $u_{i},v_{t}$ that collectively characterize the signal of each entry $(i,t)$ of the matrix [37, 20, 4, 16, 23]. Such a factor model is analogous to the low rank assumptions commonly imposed in matrix completion [19]. The second assumption specifies how the missing pattern $A_{i,t}$ was generated; for instance missing completely at random (MCAR) assumes that $A_{i,t}$ are independent to all other randomness present in the model 1 and that all entries have positive probability of being observed.

### v2

@@ 2.3 New variant: Auto nearest neighbors
$\mathsf{TSNN}$ is a generalization of $\mathsf{RowNN}$ and $\mathsf{ColNN}$ by setting one of the tuning parameters to zero (see Tab.˜1), whereas the idea underlying $\mathsf{DRNN}$ is fundamentally different from that of $\mathsf{TSNN}$; $\mathsf{DRNN}$ debiases a naive combination of $\mathsf{RowNN}$ and $\mathsf{ColNN}$ whereas $\mathsf{TSNN}$ simply boosts the number of measurements averaged upon, thereby gaining from lower variance. So we simply interpolate the two methods for some hyper-parameter $\alpha\in[0,1]$; see Tab.˜1. Notably the hyper-parameter $\eta$ for both $\mathsf{DRNN}$ and $\mathsf{TSNN}$ are identical when interpolated.
Suppose $\mu_{i,t}=\theta_{i,t}+\varepsilon_{i,t}$ in ˜2 where $\varepsilon_{i,t}$ are centered i.i.d. sub-Gaussian distributions across $i$ and $t$. When $\sigma$ is large in magnitude, $\mathsf{TSNN}$ denoises the estimate by averaging over more samples, hence providing a superior performance compared to $\mathsf{DRNN}$ in a noisy scenario. When $\sigma$ is small so that bias of nearest neighbor is more prominent, $\mathsf{DRNN}$ effectively debiases the estimate so as to provide a superior performance compared to $\mathsf{TSNN}$. The linear interpolator $\mathsf{AutoNN}$ automatically adjusts to the underlying noise level and debiases or denoises accordingly; such property is critical when applying nearest neighbors to real world data set where the noise level is unknown. We refer to Fig.˜1 for visual evidence.
@@
Figure 1: Error scaling for certain NN variants in synthetic experiments. See Sec.˜D.1 for details on the data-generating process and how the signal-to-noise ratio (SNR) is defined. Each point corresponds to the mean absolute error ± 1 standard error across 30 trials.
@@ 3 $\mathbf{N}^{2}$ Package and Interface
We now present our unified Python package, $\mathbf{N}^{2}$, for nearest neighbor algorithms for matrix completion. In particular, we provide a class structure which abstracts the estimation procedure utilized in each different nearest neighbor method as well as the the Distance and Average modules described above in Sec.˜2. On top of that, our library facilitates easy extensions to other nearest neighbors algorithms and other data types on top of scalars and distributions. For example, as long as a distance and average notion are well defined, our library can be easily applied to a matrix of images or text strings.
@@
EstimationMethod classes contain the logic to impute a missing entry such as how to use calculated distances. In the context of the modules in Sec.˜2, this class determines the weighting function or, in the context of $\mathsf{DRNN}$ and $\mathsf{AutoNN}$, how to compose estimates. We separate this from the DataType abstraction because several estimation methods can be used for multiple data types. For example, RowRowEstimator implements the $\mathsf{RowNN}$ procedure for any data type given to it, such as scalars or distributions.
@@
To use our library, a user simply has to instantiate a composite class NearestNeighborImputer with their EstimationMethod and DataType of choice. We provide constructor functions to automatically create popular NearestNeighborImputer classes such as a two-sided nearest neighbor estimator with the scalar data type. From a design pattern point of view, this is known as a Composite design pattern (gamma1993design, pg. 163). We use this design pattern so that anyone looking to customize the estimation procedure can do so for any kind of data type simultaneously. Similarly, with the exception of doubly robust estimators, each estimation procedure works out of the box with any data type that implements the DataType abstract class. The Doubly robust estimation method does not work out of the box with distributions because a subtraction operation is not well defined in the distribution space.
@@ 4 $\mathbf{N}^{2}$-Bench and Results
In this section, we evaluate several nearest neighbor algorithms provided by our library, $\mathbf{N}^{2}$, on real-world data. As part of our package, we include data loaders which automatically download the necessary datasets and format them for evaluation. These datasets and loaders comprise our proposed benchmark for nearest neighbor matrix completion algorithms, $\mathbf{N}^{2}$-Bench. We also test several existing popular matrix completion techniques (hastie2015matrix; chatterjee2015matrix). For details on our experimental setup, computing hardware, and boxplot generation, see App.˜D.
@@ 4.1 Personalized healthcare: HeartSteps
The HeartSteps V1 study (HeartSteps study for short) is a clinical trial designed to measure the efficacy of the HeartSteps mobile application for encouraging non-sedentary activity klasnja2019efficacy. The HeartSteps V1 data and its subsequent extensions have been widely used for benchmarking a variety of tasks including counterfactual inference of treatment effect dwivedi2022counterfactual; choi2024learning, reinforcement learning for intervention selection liao2019rlheartsteps, and micro-randomized trial design qian2022microrandomized. In the HeartSteps study, $N=37$ participants were under a 6-week period micro-randomized trial, where they were provided with a mobile application and an activity tracker. Participants independently received a notification with probability $p=0.6$ for $5$ pre-determined decision points per day for 40 days ($T=200$). We denote observed entries $Z_{i,t}$ as the mean participant step count for one hour after a notification was sent and unobserved entries as the unknown step count for decision points where no notification was sent. Our task is to estimate the counterfactual outcomes: the participant’s step count should they have received a different treatment (notification or no notification) than they did at specific time points during the study.
@@
In Fig.˜2(a), we compare the absolute error of the imputed values across the nearest neighbor and baseline methods. The scalar nearest neighbor methods far out-perform USVT and are on par with SoftImpute. The two distributional nearest neighbor methods far outperform all methods operating in the scalar setting; it suggests that matching by distributions collect more homogeneous neighbors, thereby decreases the bias of the method, compared to matching only the first moments as done in most scalar matrix nearest neighbor methods.
In Fig.˜2 panel (b), we show an example of an imputed entry in the distributional nearest neighbors setting. In this case, the ground truth distribution is bimodal, as the participant was largely sedentary (0 steps) with small amounts of activity. While both $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ capture the sedentary behavior of the participant, $\mathsf{KernelNN}$ is able to recover the bimodality of the original distribution whereas $\mathsf{W_{2}NN}$ cannot.
@@ 4.2 Movie recommendations: MovieLens
Figure 3: MovieLens: Estimation error for a random subsample of size 500. For experimental settings and discussion see Sec.˜4.2.
The MovieLens 1M dataset movielens contains 1 million ratings (1–5 stars) from 6,040 users on 3,952 movies. Collaborative filtering on MovieLens has long been a benchmark for matrix-completion methods: neighborhood-based algorithms sarwar2001item, latent-factor models koren2009matrix, and, more recently, nearest neighbors interpreted as blind regression under a latent–variable model li2019nearest. These assist practitioners in data-driven recommendation systems, since more accurate rating imputation directly drives better personalized suggestions and user engagement. This is a standard scalar matrix completion problem with $N=6,040$ and $T=3,952$. Each rating is an integer in $\{1,\dots,5\}$. The dataset has a very high percentage of missing values: 95.53% missing. Our task is to estimate unobserved ratings using various matrix completion algorithms. We benchmark the performance of nearest neighbors against matrix factorization by measuring absolute error on held‑out ratings. See Sec.˜D.3 for additional details on the dataset.
@@
We fit the nearest neighbor methods using a random sample of size 100 from the first 80% of the dataset to choose nearest neighbor hyperparameters via cross-validation.We then test the method on a random subsample of size 500 from the last 20% of the dataset. As observed in Fig.˜3, all nearest neighbor methods have a lower average error than USVT and a much lower standard deviation of errors, with $\mathsf{ColNN}$, $\mathsf{RowNN}$, $\mathsf{DRNN}$, and $\mathsf{AutoNN}$ performing the best out of the nearest neighbor methods. SoftImpute performs on par with the nearest neighbor methods. Note that the nearest neighbor methods perform well even while only being trained on a tiny subset of the data of size 100 out of the 1 million ratings available.
@@ 4.3 Counterfactual inference for panel data: Proposition 99
Figure 4: Nearest neighbor methods generate high-fidelity synthetic controls in counterfactual inference for panel data. For exact settings and further discussion see Sec.˜4.3.
Next we consider a panel data setting, where our goal is to estimate the effect of the California Tobacco Tax and Health Protection Act of 1988 (a.k.a. Proposition 99) on annual state-level cigarette consumption222measured as per capita cigarette sales in packs. By definition, the counterfactual cigarette consumption in California—had Proposition 99 never been enacted—is not observed. abadie2010synthetic introduce the notion of a “synthetic control” to serve as a proxy for this unobserved value based on “neighboring” control states that never instituted a tobacco tax. These states are not close in a geographical sense, but rather close due to similarities in other covariates [^3] GDP per capita, beer consumption, percent aged 15–24, and cigarette retail prices. We take a different approach and use only the observed cigarette consumption levels from the control states, of which there are 38 in total. Thus, we frame our problem as a scalar matrix completion problem with $N=39$ and $T=31$ (see ˜2). The last row in the matrix corresponds to the state of California.
@@
For each method, we use a 64-16-20 train-validation-test split and use cross validation to fit any hyperparameters. Fig.˜4 plots the various synthetic controls for California (left) and absolute error of each method on the 38 control states, for which we do observe the no-treatment values (right). From Fig.˜4(a), we see that nearest neighbor methods, in particular $\mathsf{TSNN}$ and $\mathsf{RowNN}$, are roughly on par with the gold-standard synthetic control method of abadie2010synthetic (“SC”) for estimating California’s counterfactual cigarette consumption in the post-intervention period (after 1989). This is despite the fact that the nearest neighbor methods rely on less information for the estimation task. From Fig.˜4(b), we see that all nearest neighbor methods, with the exception of $\mathsf{ColNN}$, achieve similar error levels as the synthetic control baseline. $\mathsf{RowNN}$ achieves even lower error levels. See supplementary experiment details in Sec.˜D.4.
@@ 4.4 Efficient LLM evaluation: PromptEval
The rapid advancement of LLMs have placed them at the center of many modern machine learning systems, from chatbots to aids in medical education geathers2025benchmarking. In practice, system architects want to strike the right balance of real-world performance and cost, but navigating this Pareto frontier is a daunting task. 2024 alone saw at least 10 new models from Anthropic, Google, Meta, and OpenAI, not even counting the multitude of open-source fine-tuned models built on top of these. On specific tasks, smaller, fine-tuned models may even outperform the latest frontier models, in addition to being more cost effective.
We investigate how matrix completion, specifically nearest neighbor methods, can alleviate some of these burdens. We use the PromptEval dataset polo2024efficient, which evaluates $15$ open-source language models (ranging in size from 3B to 70B parameters) and $100$ different prompting techniques across the 57 tasks of the MMLU benchmark (hendrycks2020measuring). In practice, the performance of a model depends—sometimes dramatically—on the precise input prompt. This suggests that we need to consider the performance of a model across a wide range of prompts, rather than any one prompt in particular. Thus, we model this problem as a distributional matrix completion problem with $N=15$, $T=57$, and $n=100$. Given one of 57 tasks, we aim to accurately characterize the performance of each model without resorting to exhaustive evaluation. Nearest neighbors leverage commonalities across models and tasks to estimate the performance distribution of each entry, which was otherwise not considered in polo2024efficient; previous literature achieves efficient evaluation per model and task in isolation without leveraging any across model / task information.
@@
Figure 5: Distributional nearest neighbor methods enable efficient LLM evaluation on MMLU. We estimate LLM score distributions across all models and tasks given only a limited number of model-task evaluations, determined by the propensity pp. See Sec.˜4.4 for a detailed discussion.
@@
We randomly include each entry in the matrix independently with probability $p\in\mathopen{}\mathclose{{\left\{0.3,0.5,0.7}}\right\}$ and impute the missing entries using the $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ methods of Tab.˜1. For each method, we consider both the the row-wise and column-wise variants. Fig.˜5(a) reports the mean Kolmogorov-Smirnov (KS) distance between the imputed and ground-truth distributions across the entries in the test set for varying missingness values. As expected, estimation error decreases as $p$ increases. Fig.˜5(b) visualizes the imputed distributions using row-wise $\mathsf{KernelNN}$ and column-wise $\mathsf{W_{2}NN}$ (at $p=0.7$) for a select entry, along with the ground-truth distribution. Even with 30% of matrix entries missing, distributional NN methods are able to recover the underlying distribution.
@@ Appendix A Structural assumptions
Provable guarantees of nearest neighbors in matrix settings ˜2 can be shown when structural assumptions are imposed on the distributions $\mu_{i,t}$ and the missingness $A_{i,t}$. We collect existing results from [li2019nearest, dwivedi2022doubly, choi2024learning, feitelberg2024distributional, sadhukhan2024adaptivity, sadhukhan2025adaptivelyweightednearestneighborsmatrix]. Given data with missing observations from ˜2, the practitioner is interested in learning information of the distributions, e.g., mean of the distributions $\{\theta_{i,t}=\int xd\mu_{i,t}(x)\}.$
The first assumption specifies the factor structure on the mean; that is, there exists latent factors $u_{i},v_{t}$ that collectively characterize the signal of each entry $(i,t)$ of the matrix [li2019nearest, dwivedi2022counterfactual, agarwal2023causal, choi2024learning, feitelberg2024distributional]. Such a factor model is analogous to the low rank assumptions commonly imposed in matrix completion [candes2012exact]. The second assumption specifies how the missing pattern $A_{i,t}$ was generated; for instance missing completely at random (MCAR) assumes that $A_{i,t}$ are independent to all other randomness present in the model ˜2 and that all entries have positive probability of being observed.